Dr. Bryan Carpenter

Department of Electronics and Computer Science University of Southampton Highfield Southampton S09 5NH

Tel: 0703 593081

email: dbc@uk.ac.soton.ecs

March 1, 1994

Dr. Philip Wadler, Computing Science Department, The University, Glasgow G12 8QQ

Dear Dr. Wadler,

I’m sorry that it has taken so long to referee this paper.

My general feeling is that it contains useful material, which would probably be of interest to JFP readers, but the paper has been written rather carelessly. I recommend that it could be accepted with revisions, but it may need refereeing again.

Yours sincerely,